# Learning Loss Landscapes in Preference Optimization

### v1

@@ Abstract
We present an empirical study investigating how specific properties of preference datasets, such as mixed-quality or noisy data, affect the performance of Preference Optimization (PO) algorithms. Our experiments, conducted in MuJoCo environments, reveal several scenarios where state-of-the-art PO methods experience significant drops in performance. To address this issue, we introduce a novel PO framework based on mirror descent, which can recover existing methods like Direct Preference Optimization (DPO) and Odds-Ratio Preference Optimization (ORPO) for specific choices of the mirror map. Within this framework, we employ evolutionary strategies to discover new loss functions capable of handling the identified problematic scenarios. These new loss functions lead to significant performance improvements over DPO and ORPO across several tasks. Additionally, we demonstrate the generalization capability of our approach by applying the discovered loss functions to fine-tuning large language models using mixed-quality data, where they outperform ORPO.
@@ 1 Introduction
Learning from human feedback is a paradigm that enables the alignment of complex agents to human preferences, and has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023). In particular, fine-tuning pretrained LLMs with human preferences has become a popular strategy to adapt them to specific tasks and to improve their safety and helpfulness.
Most LLM alignment pipelines begin with a supervised fine-tuning (SFT) step, which involves supervised next-token prediction on a dataset of high-quality responses and leads to a reference policy. The reference policy is further optimized using the human preference data, typically through either Reinforcement Learning from Human Feedback (RLHF)  (Christiano et al., 2017) or Direct Preference Optimization (DPO)  (Rafailov et al., 2024), or one of their several variants. RLHF consists of learning a reward model consistent with human preferences and then using Reinforcement Learning (RL) techniques such as REINFORCE (Sutton et al., 1999) and Proximal Policy Optimisation (PPO)  (Schulman et al., 2017) to maximize the total expected reward. In contrast, DPO and its variations, e.g. odds ratio preference optimization (ORPO) (Hong et al., 2024), bypass explicit reward models entirely and optimize directly on preference data, implicitly learning the reward.
While RL-based methods offer stronger theoretical guarantees and often lead to higher performance (Song et al., 2024; Xu et al., 2024), offline approaches such as DPO have gained traction due to their simplicity and the ability to leverage preexisting high-quality datasets. In contrast to PPO, where data collection and labeling are performed iteratively after each update, DPO and its modifications allow for more efficient training, avoiding the high computational costs, need of additional sample labelling and complexity of RL methods. Specifically, PPO requires careful parameter tuning (Yuan et al., 2023) and involves simultaneous training of multiple models (the reward model, language model, and critic), which can be prohibitive in most hardware setups.
@@
The performance of offline algorithms such as DPO, particularly on noisy or low-quality datasets, has been a subject of debate (Chowdhury et al., 2024), with limited empirical results available. In this work, we provide a comprehensive analysis of PO algorithms, examining their behavior on automatically generated preference datasets. We perform this analysis in MuJoCo environments (Todorov et al., 2012), where the underlying reward structure is well defined and offers a clear performance metric to compare agents. In particular, we focus on ORPO, as we have found that it largely outperforms DPO in all settings we have considered. Our findings indicate that ORPO exhibits distinct failure modes when applied to specific low-quality or noisy datasets. These failure modes are present in practical LLM applications, raising concerns about using mixed-quality datasets for PO.
To find algorithms that are capable of dealing with these more difficult settings, we introduce a novel framework for PO based on mirror descent (Nemirovski & Yudin, 1983), which generalizes DPO and ORPO and allows to search for new Preference Optimization (PO) algorithms. We perform this search over the space of mirror maps using evolutionary strategies (ES), as they have been shown to be effective for discovering algorithms (Lu et al., 2022; Jackson et al., 2024). We outline our contributions in detail below.
@@
    We perform a systematic analysis of ORPO on preference datasets with varying levels of data quality, noise levels, initial policy, and judge temperature.
@@
    We introduce a novel family of offline PO algorithms using mirror descent, which can be easily parameterized and which we explore via ES.
@@
    For each failure mode we identify, we find and describe an algorithm within our framework that outperforms ORPO. Additionally, we show that allowing losses to vary based on the percentage of training progress drastically boosts performance in some settings.
@@
    We demonstrate that PO algorithms discovered on MuJoCo can generalize to LLM tasks. In particular, for a highlighted failure setting, we show that our discovered algorithm improves performance w.r.t. the ORPO baseline. This finding showcases the efficacy of analyzing PO algorithms in simpler, less computationally expensive environments, as it is possible to obtain insights relevant for LLM applications.
@@ 2 Preliminaries
Let $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,T,\mu)$ denote an episodic Markov Decision Process, where $\mathcal{S}$ and $\mathcal{A}$ are respectively the state and action spaces, $P(s^{\prime}\mid s,a)$ is the transition probability from state $s$ to $s^{\prime}$ when taking action $a$, $r(s,a)\in[0,1]$ is the reward function, $T$ is the maximum episode length, and $\mu$ is a starting state distribution. A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$. Let $\tau=\{(s_{t},a_{t})\}_{t=0}^{T-1}$ denote a trajectory of length $T$ and, with a slight overload of notation, let $\pi(\tau)=\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})$ and $r(\tau)=\sum_{t=0}^{T-1}r(s_{t},a_{t})$. Lastly, let $\pi(\cdot|\tau)$ be the product distribution of $\pi(\cdot|s_{0}),\dots,\pi(\cdot|s_{N-1})$.
Let $\mathcal{D}=\{(s_{0}^{i},\tau_{w}^{i},\tau_{l}^{i})_{i=1}^{N}\}$ be a preference dataset, where each tuple $(s_{0},\tau_{w},\tau_{l})$ consists of a starting state $s_{0}$ and two trajectories with starting state $s_{0}$. Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$. We assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
$\mathbb{P}(\tau_{w}\succ\tau_{l})=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$ (1)
@@
where $\sigma$ is the sigmoid function and $\eta$ is a temperature parameter. Our objective is to exploit the dataset $\mathcal{D}$ to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward of an episode, that is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau):=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\sum_{t=0}^{T-1}r(s_{t},a_{t}).$ (2)
We consider an offline setting, where we do not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
2.1 Preference optimization
There are several methods in the literature to optimize the objective in (2) using a preference dataset $\mathcal{D}$. We provide here a short review of the main pipelines (Ouyang et al., 2022). Two common independent preliminary steps are reward modelling (RM) and supervised fine tuning (SFT). RM aims to obtain an estimate of the true reward function, and is usually framed as a maximum likelihood estimation problem for a Bradley-Terry preference model, i.e. find
$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}\log\sigma(r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l})/\eta),$ (3)
@@ 2.1 Preference optimization
for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$ and for $\eta\geq 0$. SFT is an initial alignment phase, where the starting policy $\pi_{0}$ is trained to imitate high-quality demonstration data. In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$, typically doing one epoch over the dataset. We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure. The new objective we want to optimize is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$ (4)
where the $D_{\mathtt{KL}}$ represents the $\mathtt{KL}$-divergence and is introduced to prevent the policy from moving too far away from the dataset distribution. The expressions in (3) and (4) can be optimized sequentially, first obtaining a reward estimate in (3) and then optimizing (4) with PPO using the reward estimate in place of the true reward. Alternatively, the optimization problems in (3) and (4) can be solved implicitly using DPO or ORPO. We proceed to discuss each of these methods.
PPO
PPO is recognized as one of the preferred methods for optimizing (4) when the necessary computing resources are available, as demonstrated by its success in training state of the art models like GPT-4 (Achiam et al., 2023) and Claude (Antropic, 2023). However, it presents a complex pipeline where one needs to effectively train a reward model, perform SFT and then optimize (4), where each phase has a different set of hyper-parameters. Additionally, storing the reward model, the reference agent and the current agent in memory is impractical in most setups and often requires sacrificing other aspects, such as batch-size. Besides it computational costs, PPO is known to be prone to reward overoptimization (Coste et al., 2024).
@@
Direct Preference Optimization (DPO) bypasses the need for an explicit reward model by using the agent itself to implicitly represent the reward model. It consists in optimizing the objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$ (5)
which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3). Refer to Appendix A for details. Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs as an alternative to PPO (Yuan et al., 2024; Jiang et al., 2024).
A known issue of DPO is that it pushes probability mass away from the preference dataset and to unseen responses, which can cause the final policy to deviate significantly from the reference policy, even when the reference policy aligns well with human preferences. In contrast, PPO can leverage the generalization capabilities of the (learned) reward model to generate responses beyond the preference dataset distribution, while the $\mathtt{KL}$-divergence penalty can provide additional regularization. To mitigate the risks described above, DPO is usually only applied for a few epochs.
@@
ORPO is a more recent algorithm that aims to further simplify the training pipeline and, concurrently, to address the distribution shift issue present in DPO. ORPO merges the SFT and DPO steps into one, which optimizes the unified objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\right]$ (6)
where $\operatorname{odds}_{\pi}(\tau)=\pi(\tau)/(1-\pi(\tau))$. ORPO gets rid of the need for a reference model by adding an SFT term to the preference optimization objective function, which substitutes the role of the SFT step and of the reference model in preventing the optimized policy from moving too far away from the dataset distribution. Additionally, the SFT term prevents pushing probability mass away from the preference dataset, addressing the distribution shift issue present in DPO.

### v2

# Meta-Learning Objectives for Preference Optimization

@@ Abstract
Evaluating preference optimization (PO) algorithms on LLM alignment is a challenging task that presents prohibitive costs, noise, and several variables like model size and hyper-parameters. In this work, we show that it is possible to gain insights on the efficacy of PO algorithm on much simpler benchmarks. We design a diagnostic suite of MuJoCo tasks and datasets, which we use to systematically evaluate PO algorithms, establishing a more controlled and cheaper benchmark. We then propose a novel family of PO algorithms based on mirror descent, which we call Mirror Preference Optimization (MPO). Through evolutionary strategies, we search this class to discover algorithms specialized to specific properties of preference datasets, such as mixed-quality or noisy data. We demonstrate that our discovered PO algorithms outperform all known algorithms in the targeted MuJoCo settings. Finally, based on the insights gained from our MuJoCo experiments, we design a novel PO algorithm that significantly outperforms existing baselines in an LLM alignment task.
@@ 1 Introduction
Learning from human preferences (Christiano et al., 2017) is a paradigm which enables the alignment of machine learning systems to relative human preferences, without requiring access to absolute rewards. While the framework was developed for robotic and games applications with experiments on MuJoCo simulations and Atari (Akrour et al., 2012; Biyik & Sadigh, 2018; Ibarz et al., 2018), this paradigm has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023). In particular, fine-tuning pre-trained LLMs with human preferences has become a popular strategy to adapt them to specific tasks and to improve their safety and helpfulness.
Within this framework, Reinforcement Learning from Human Feedback (RLHF) is one of the most popular methods. It consists in learning a reward function using a preference dataset and then optimizing the estimated reward using Reinforcement Learning methods such as Proximal Policy Optimization (PPO)  (Schulman et al., 2017). However, the training pipeline of RLHF is quite complex, which is why implicit approaches such as Direct Preference Optimisation (DPO)  (Schulman et al., 2017) have gained traction thanks to their simplicity. These methods do not learn a reward model but estimate it implicitly using the policy of the agent. Many follow ups to DPO have been proposed  (Yuan et al., 2023; Zhao et al., 2023; Azar et al., 2024; Xu et al., 2024a; Hong et al., 2024; Park et al., 2024; Meng et al., 2024), but comparing their performance in LLM alignment is a complex task that incurs high costs, noise, and the inherent difficulty in judging a response better than another.
@@
In this work, we provide a comprehensive analysis of PO algorithms, examining their behavior on automatically generated preference datasets. We return to the roots of RLHF by performing this analysis in MuJoCo environments and datasets, where the underlying ground-truth reward structure is well defined and offers a clear performance metric to compare agents. In particular, we design a task where a pre-trained agent has to adhere to a new stylistic constraint, emulating the typical conditions of LLM fine-tuning. Our findings indicate that many PO algorithms present distinct failure modes when applied to specific mixed-quality or noisy datasets.
Moreover, we introduce a framework for finding PO algorithms. Specifically, we define a class of PO algorithms based on mirror descent (Nemirovski & Yudin, 1983), which generalizes DPO and ORPO for particular choices of the mirror map. We then show that this class can be easily parametrized and searched using evolutionary strategies (ES), optimizing for the final performance of the trained policy, as measured by the ground truth reward.
For each setting we consider, we discover an algorithm that significantly outperforms all baselines. Analyzing the discovered algorithms, we find that the main difference between them and the baselines is that they keep optimizing the policy of the agent well after the probability of generating the chosen trajectory has surpassed the probability of generating the rejected one. We use this insight to design a new PO function, Constrained Linear Policy Optimization (CoLPO) that demonstrates promising results in an LLM alingment task. We summarize our contributions below.
@@
    We perform a systematic evaluation of eight of the most popular PO algorithms on automatically generated preference datasets with varying levels of data quality, noise levels and initial policy. We see that most existing algorithms struggle when dealing with noise and mixed-quality date.
@@
    We introduce a novel family of offline PO algorithms using mirror descent, named Mirror Preference Optimization (MPO), which can be easily parameterized and explored via ES.
@@
    For both noisy and mixed-quality settings, we find and describe a PO algorithm within our framework that largely outperforms all the considered baselines in our MuJoCo benchmark.
@@
    We demonstrate that takeaways from our analysis on the MuJoCo setting, as well as the characteristics of the discovered PO algorithms, can be successfully transferred onto LLM tasks. In particular, we show that our CoLPO algorithm significantly improves upon the ORPO, CPO, and SLiC-HF baselines.
Overall, we show that it is possible to gain relevant insights for the design of LLM alignment algorithms through a systematic analysis in simple, computationally inexpensive environments such as MuJoCo.
@@ 2 Preliminaries
A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$. Let $\tau=\{(s_{t},a_{t})\}_{t=0}^{T-1}$ denote a trajectory of length $T$ and, with a slight overload of notation, let $\pi(\tau)=\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})$ and $r(\tau)=\sum_{t=0}^{T-1}r(s_{t},a_{t})$. Lastly, let $\pi(\cdot|\tau)$ be a distribution over $(\Delta(\mathcal{A}))^{T}$ defined as $\pi(\cdot|s_{0})\times\dots\times\pi(\cdot|s_{N-1})$.
@@
Our objective is to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward of an episode, that is
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau)$ (1)
$:=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t},s_{t+1}}\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
where $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
Let $\mathcal{D}=\{(s_{0}^{i},\tau_{w}^{i},\tau_{l}^{i})_{i=1}^{N}\}$ be a preference dataset, where each tuple $(s_{0},\tau_{w},\tau_{l})$ consists of a starting state $s_{0}$ and two trajectories with starting state $s_{0}$. Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$. Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
$\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}$ (2)
$=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$
where $\sigma$ is the sigmoid function and $\eta$ is a temperature parameter. In this work, we consider an offline training setting, where the agent aim to solve the optimization problem in (1) but only has access to the the dataset $\mathcal{D}$ and cannot collect further data. We also assume the agent does not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
2.1 Alignment to preference feedback
There are several algorithms in the literature to optimize the objective in (1) using a preference dataset $\mathcal{D}$. We describe supervised fine-tuning (SFT), DPO and ORPO, as they are among the most popular and as many methods can be seen as a variation of one of these algorithms.
SFT
@@ 2.1 Alignment to preference feedback
SFT is an initial alignment phase, where the policy $\pi_{0}$ is trained to imitate high-quality demonstration data. In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$. We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure.
@@
Direct Preference Optimization (DPO) consists in solving a maximum likelihood estimation problem and a policy optimization problem in a single step. The maximum likelihood estimation problem is the one to find an estimate of the true reward function that governs how the preferences are expressed, that is
$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}g\circ\sigma((r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l}))/\eta),$ (3)
for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$, for an increasing function $g:\mathbb{R}\leftrightarrow\mathbb{R}$, and for $\eta\geq 0$. Here, typically, $g$ is set to be the logarithm function. The policy optimization problem is the one to maximize the expected reward, that is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (4)
$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\widehat{r}(s_{t},a)-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$
where $D_{\mathtt{KL}}$ represents the $\mathtt{KL}$-divergence and is introduced to prevent the policy from moving too far away from the dataset distribution.
DPO merges these two problems by using the agent itself to implicitly represent the reward model. It consists in optimizing the objective
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$ (5)
$\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$
which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3), with $g=\log$. Refer to Appendix C for details. Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs (Yuan et al., 2024; Jiang et al., 2024).
A known issue of DPO is that it pushes probability mass away from the preference dataset and to unseen responses (Xu et al., 2024b), which can cause the final policy to deviate significantly from the reference policy, even when the reference policy aligns well with human preferences. To mitigate this risk, DPO is usually applied for a few epochs.
@@
ORPO further simplifies the training pipeline and addresses the distribution shift issue present in DPO. It merges the SFT and DPO steps into one, optimizing the unified objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\bigg{[}\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}$ (6)
$+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\bigg{]}$
where $\operatorname{odds}_{\pi}(\tau)=\pi(\tau)/(1-\pi(\tau))$. ORPO gets rid of the need for a reference model by adding an SFT term to the preference optimization objective function, and uses this term to prevent the optimized policy from moving too far away from the dataset distribution. Additionally, the SFT term prevents pushing probability mass away from the preference dataset, addressing the distribution shift issue present in DPO.
Research on preference optimization has been very active and many methods have been proposed. We present a summary of the most popular algorithms in Table 1 and a brief discussion in Appendix A.
Beyond these implicit algorithms, there are several other methods that explicitly solve the maximum likelihood problem in (3) and use the learned reward model to optimize the objective in (4) with an RL algorithm. Overall, RLHF is a superior approach and the industry standard, but is more computationally expensive and complex to implement. For a detailed discussion and comparison between DPO-like methods and PPO, refer to Appendix F.
Method Objective
RRHF (Yuan et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max\left(0,-\frac{1}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)+\frac{1}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)\right)$
SLiC-HF (Zhao et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max(0,\delta-\log\pi_{\theta}(\tau_{w}|x)+\log\pi_{\theta}(\tau_{l}|x))$
DPO (Rafailov et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}\right)$
IPO (Azar et al., 2024) $-\left(\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}-\frac{1}{2\tau}\right)^{2}$
CPO (Xu et al., 2024a) $\log\pi_{\theta}(\tau_{w}|x)+\log\sigma(\beta\log\pi_{\theta}(\tau_{w}|x)-\beta\log\pi_{\theta}(\tau_{l}|x))$
ORPO (Hong et al., 2024) $\log\pi(\tau_{w})+\lambda\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)$
R-DPO (Park et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}+(\alpha|\tau_{w}|-\alpha|\tau_{l}|)\right)$
SimPO (Meng et al., 2024) $\log\sigma\left(\frac{\beta}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)-\frac{\beta}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)-\gamma\right)$
Table 1: A summary of popular PO algorithms. The objective in each row is to be maximized and $(\tau_{w},\tau_{l})\sim\mathcal{D}$.